# Sparse Mean Field Load Balancing in Large Localized Queueing Systems

### 20 Dec 2023 v1

@@ 3 Experiment Results
Comparison to SAC
@@
To begin, while training for SAC, we observed that on increasing the number of agents, the convergence to a locally optimal policy takes longer (using one core of Intel Xeon Gold 6130), making it not too feasible for the larger setups we consider in this work. See Figure 4(a) for time taken to learn a SAC policy on a $1$-D cyclic topology with $\Delta t=3$, $N=M$, and same computational resources for all training setups of $N$. Our implementation of SAC was adapted from Li (2021). Furthermore, Figure 4(b) shows the performance of the learned SAC policy as compared to other algorithms. Although performance did improve as the number of agents rises, indicating the scalability of SAC to many agents, the performance of SAC remained suboptimal. Due to these limitations, we did not investigate the SAC algorithm further in our experiments.
Finite system performance
The second result of interest is the performance analysis on the $1$-D cyclic graph over a range of $\Delta t$ for the learned MF-R policy as compared to the NA-PS, JSQ, RND and OWN algorithms, see Figure 5. The number of agents (queues) used to generate the graph are $N\in\{9,21,91,901,3501,5001\}$, with every agent always having the fixed degree $d=2$. For this graph, we also trained an NA-PS policy on a sufficiently small size with $N=M=5$ to obtain convergence even with batch and minibatch sizes of $50000$ and $8000$. We used PPO with parameter sharing so that the learned policy could be evaluated on any graph size.
It can be seen that JSQ is the best strategy at $\Delta t=1$, while the performance of MF -R is very close to JSQ. And as expected, the performance of JSQ deteriorates with increasing synchronisation delay $\Delta t$. For an intermediate range of $\Delta t\in\{3,5,7\}$, neither JSQ nor RND are the optimal load balancing policies. Here, our learned MF -R policy performs best and approaches the optimal RND performance as $\Delta t$ increases to $10$. As discussed earlier, it has already been proven in certain scenarios that as $\Delta t\to\infty$, RND is the optimal load balancing policy Mitzenmacher (2001). The learned NA-PS policy has a similar trend to our MF -R policy, but is unable to outperform it even for small systems. For the CYC -1D, CCC, and TORUS graphs with homogeneous degrees, the performance of OWN in the limit – sending only to its own queue – is equivalent to RND, since for homogeneous degrees, sending to your own queue results in the same packet arrival rate at each queue as randomly sending to any queue in your neighbourhood. In Figure6, it can also be seen that as the size of the graph increases, the performance of the policy converges and approaches the learned MF -R policy, confirming our theoretical mean field analysis. The convergence also confirms that the $N$ chosen for training was large enough to represent the mean field system.
Figure 9: (a): The training for TORUS with and without RNN policies converges almost to the same return. (b): Evaluation of the learned policy using different observations. MF-R uses only own queue state, MF-N additionally uses neighbour queue states, and MF-G uses state of all queues in the system.
Performance in large systems
In Figure 7 each subfigure shows results for a different topology. We show the performance over a range of $\Delta t\in\{1,2,\ldots,10\}$ for large sparse networks. Figure 7(a) is a compact representation of Figure 5 over all $\Delta t$ and $N=5001$. Figure 7(b) is for the CCC graph with $d=3$ and $N=4608$. Figure 7(c) shows the performance for the TORUS graph with $d=4$, $o_{t}=70$ and $N=4900$, while Figure 7(d) is for the CM graph for $N=5001$ with uniformly distributed degrees in the set $d\in\{2,3\}$. Our learned MF-R policy is compared to JSQ, RND and OWN algorithms. The following observations can be made for all the subfigures: (i) For small $\Delta t=1$ JSQ performs better than all the algorithms which is theoretically guaranteed, however performance of MF-R is very close to it; (ii) For in between range of $\Delta t=\{2,3,\ldots,7\}$ MF-R has the best performance in terms of packets dropped; (iii) Except in the CM graph (where the varying degrees lead to differing packet arrival rates between OWN and RND), we find that OWN and RND again have equivalent performance by regularity of the graph, as discussed in the previous paragraph; (iv) For higher $\Delta t\in\{8,9,10\}$ our MF-R policy performance coincides with the RND policy, where the RND policy has been theoretically proven to be optimal as the synchronization delay $\Delta t\to\infty$ Mitzenmacher (2001). Hence, we find that our MF-R policy performs consistently well in various topologies.
Ablation on topological assumptions
As mentioned in Section2.1, the theoretical analysis and work in Lacker et al. (2023) does not apply to the Bethe topology, and the effect is further supported by our experimental results. Figure8 shows the performance comparison for Bethe lattice graphs of the learned MF-R policy with JSQ, RND, and OWN algorithms over different $\Delta t$. We show the result only for a large Bethe network, with $o_{b}=11$, $d=3$ and $N=6142$. OWN outperforms RND more significantly than in CM, since disproportionately many packets arrive at nodes connected to leaf nodes. It can be seen that the system behaves as expected when $\Delta t$ is small, with JSQ quickly outperformed by MF-R, while MF -R eventually performs worse than OWN as $\Delta t$ increases. This behaviour is due to the fact that in structures such as regular trees, the central node as the root node is not sufficient to represent the performance of all nodes, especially the many leaf nodes that behave differently even in the limit. Because of this loss of regularity, theoretical guarantees fail when scaling large Bethe or similar topologies, as explained in Section2.1. Nevertheless, we see that the learned policy can outperform existing solutions in certain regimes, e.g., at $\Delta t=3$.
RNNs and decentralized execution
We additionally trained our MF -R policy with recurrent neural networks in PPO, as is typical for partially-observed problems Ni et al. (2022), MF-RNN. In Figure 9(a), we see that their effect is negligible and can even be negative. Finally, we used partially observed decentralized alternatives to the empirical distribution as input to the learned MF-R policy in the evaluation, see Figure 9(b). In MF-N we use the distribution of neighbours’ queue states, in MF-G we use the empirical distribution of all queues in the system, and in MF -R only the agent’s own queue state information is used as a one-hot vector. Similar performances indicate that the learned policy can be executed locally without using global information.
@@
Figure 10: (a) Performance comparison for increased buffer capacity of each queue to $20$. The system utilization was increased to ensure occurrence of packet drops in a limited amount of time. (b) Performance comparison of MF-R when the servers are considered to be heterogeneous. Additional comparison was done with the Shortest-expected-delay algorithm Selen et al. (2016), which is the state-of-the-art when the servers are heterogeneous.
Increased buffer size
@@
We also conducted experiments where we considered the servers to be heterogeneous with randomly assigned speed of fast(rate $2$) or slow(rate $1$). The workload was the same; modulating between $[0.9,0.6]$. Fig. 10(b) shows that our algorithm has comparable performance to RND and OWN while outperforming JSQ. We also compared with Shortest-expected-delay (SED), the state-of-the-art load balancing algorithm for heterogeneous servers Selen et al. (2016), which performs better since it makes decision based on both the neighbors’ queue state information and the server speeds. We believe our model is not an ideal fit for this setting and further work needs to be done in this direction.
4 Conclusion
Overall, we have learned efficient algorithms for load balancing in highly local queuing systems with some degree of theoretical guarantees through sparse mean-field approximations. The approach has been positively evaluated in comparison to well-known baselines, and can scale to queueing systems of arbitrary size with arbitrary synchronization delays. Future work could attempt to quantify rates of convergence for approximate optimality guarantees of MFC policies under the framework of Lacker et al. (2023), or consider dynamic programming principles. One could also look into the analysis of graph structures such as Bethe and regular trees, for which the current sparse graph mean-field modelling does not suffice.
Appendix A
Proof of Theorem 1.
We apply the framework of Lacker et al. (2023), which does not include actions or rewards, by including actions and rewards via separate time steps. Specifically, each time step is split in three, and we define the agent state space $\mathcal{X}\coloneqq\mathcal{Z}\cup(\mathcal{Z}\times\mathcal{A})\cup(\mathcal{Z}\times\mathcal{A}\times[0,D_{\mathrm{max}}])$ for each of the agents, indicating at any third decision epoch $t=0,3,6,\ldots$ the state of its queue. After each such epoch, the agent states will contain the state of its queue together with its choice of action at times $t=1,4,\ldots$, and lastly also the number of packet drops at times $t=2,5,\ldots$. Here, $D_{\mathrm{max}}$ is the maximum expected number of packet drops and is given by the $(d+1)$ times the maximum per-scheduler arrival rate $\lambda_{\mathrm{max}}=\lambda_{h}$.
We formally rewrite the system defined earlier, using the symbol $\mathbf{x}$ for the state of the rewritten system instead of $\mathbf{z}$. As a result, each agent $i$ is endowed with a random local $\mathcal{X}$-valued state variable $X^{G,x}_{i}(t)$ at each time $t$. For the dynamics, we let $S^{\sqcup}(\mathcal{X})$ denote the space of unordered terminating sequences (up to some maximum degree) with the discrete topology as in Lacker et al. (2023), and define in the following a system dynamics function $F\colon\mathcal{X}\times S^{\sqcup}(\mathcal{X})\times\Xi\to\mathcal{X}$ returning a new state for any current local state $X^{G,x}_{i}(t)$, any current $S^{\sqcup}(\mathcal{X})$-valued neighbourhood queue fillings $X^{G,x}_{N_{i}}(t)$ and i.i.d. sampled $\Xi$-valued noise $\xi_{i}(t+1)$. More precisely, we use
$X^{G,x}_{i}(t+1)=F[X^{G,x}_{i}(t),X^{G,x}_{N_{i}}(t),\xi_{i}(t+1)]$
For the first step, fixing any $\pi$, we let $\xi_{i,z}\sim\mathrm{Bernoulli}(\pi_{\lfloor t/3\rfloor}(z))$, for all $z\in\mathcal{Z}$, to sample action from $\pi$. Deferring for a moment the second step, for the new state in the third step, we use the Kolmogorov forward equation for the queue state during the epoch for any initial $(x,\mu)$: $P(x,\mu)=\exp{\mathbf{Q}(x,\mu)}\cdot e_{x_{1}}\in\Delta^{\mathcal{Z}}$ with unit vector $e_{x_{1}}$, which results from Poisson thinning, as the equivalent arrival rate at a queue of an agent currently in state $x$ with neighbourhood $\mu$ will be given by $\lambda\left(1-x_{2}+\frac{\mu(\mathbf{1}_{\mathcal{Z}\times\{1\}\times S^{\sqcup}(\mathcal{X})})}{d}\right)$. Thus, we have the transition rate matrix $\mathbf{Q}(x,\mu)$ with $\mathbf{Q}_{i,i-1}=\lambda\left(1-x_{2}+\frac{\mu(\mathbf{1}_{\mathcal{Z}\times\{1\}\times S^{\sqcup}(\mathcal{X})})}{d}\right)$, $\mathbf{Q}_{i-1,i}=\alpha$ for $i=1,\ldots,B$, and where $\mathbf{Q}_{i,i}=-\sum_{j}\mathbf{Q}(\lambda,z)_{j,i}$ for $i=0,\ldots,B$, and zero otherwise. Therefore, the next queue filling is sampled as $\xi_{i,x,\mu}\sim\mathrm{Categorical}(P(x,\mu))$. Lastly, for the second step we consider the non-random conditional expectation of packet losses during any epoch: $\xi_{i,x_{1},x_{2},\mu}=[\exp{\bar{\mathbf{Q}}(x,\mu)}\cdot e_{x_{1}}]_{B+1}$ under the executed actions (allocations), analogously to the new state, by adding another row to the transition matrix $\mathbf{Q}(x,\mu)$, giving $\bar{\mathbf{Q}}(x,\mu)$ where $\bar{\mathbf{Q}}_{B+1,B}=\lambda\left(1-x_{2}+\frac{\mu(\mathbf{1}_{\mathcal{Z}\times\{1\}})}{d}\right)$, i.e. counting all the expected packet arrivals whenever the queue is already full ($B$).
$\mu^{G_{n},\mathbf{z}_{n}}(t)\to\mathcal{L}(X^{G,x}_{\varnothing}(t))\quad\text{in distribution in $\mathcal{P}(\mathcal{X})$}$ (3)
since we split any time step into three by the prequel, the continuous mapping theorem, and dominated convergence, where we use the continuous function $r^{\infty}(\mu)=\int\mathbf{x}_{3}1_{x\in\mathcal{Z}\times\mathcal{A}\times[0,D_{\mathrm{max}}]}\mu(\mathrm{d}x)$, where $\mu=\mathcal{P}(\mathcal{X})$, to sum up the expected packet drops in the system, since the integrand is continuous and bounded (under the sum topology for the union in $\mathcal{X}$, and the product topology for products in $\mathcal{X}$). ∎
Proof of Corollary 1.1.
Define the optimality gap: $\Delta J\coloneqq J^{G,\mathbf{z}}(\pi_{i})-\max_{j\neq i}J^{G,\mathbf{z}}(\pi_{j}).$ Then, by Theorem 1, there exists $n$ such that: $\max_{i}\left|J^{G_{n},\mathbf{z}_{n}}(\pi_{i})-J^{G,\mathbf{z}}(\pi_{i})\right|\leq\frac{\Delta J}{2}.$ Therefore,
is the desired conclusion. ∎

### 22 Mar 2024 v2

@@ D.1. Simulator analysis
Figure 7. MF-R is with symmetric actions, MF-L is learning actions for each neighbor separately but without knowing the exact state of the neighbor, MF-NGH is learning the action for each possible non-repeating combination of states for agents’ own state and states of the neighbors. This experiment is done for the CYC-1D graph.
D.2. Different policies
For the sake of completeness, we also compared different policies for the setup of CYC-1D graph at $\Delta t=1$. First is the MF-R where the policy learned tells you the probability with which to send to your queue or randomly to one of your neighbors. Second is the MF-L where the policy learned tells you the probability with which to send to your queue or ones of your neighbors. Lastly, is the MF-NGH where a separate action is learned for every possible combination of your state and state of your neighbors. However, the neighbors are still kept anonymized by not learning a separate policy for neighbors’ state $[3,5]$ and $[5,3]$ for $d=2$, meaning that it does not matter which of your neighbors has queue filling $3$ or $5$, a single policy is learned for these combinations. In Fig. 7 it can be seen that MF-R performs the best, and we believe this is because it has the smallest action space, so learning an optimal policy is more probable. Note that the input observation was the same for all, which is the empirical distribution of the agents’ state as explained in Section 2.4.1. Hence, using any kind of action space is feasible, one just needs to consider the increase in training time and resources when the action space increases. Based on this we have used the MF-R model and policy for all our experiments.
Figure 8. (a): SAC training time to convergence increases with the number of agents, making it difficult to train and use for larger setups. (b): SAC performs worse than our proposed MF-R policy for a number of agents between $5$ and $101$.
D.3. Comparison to SAC
@@
The scalable-actor-critic (SAC) (Lin et al., 2021) method learns a discrete policy of sending to any one of the $d$ accessible queues. Each agent learns an individual policy while using as observation the agent’s own queue state and also the queue states of all its neighbours. However, a trained policy cannot be used for an arbitrary number of agents since the policy of an agent is influenced by its neighbours states as well, which is not assumed to be the same for every agent. To begin, while training for SAC, we observed that on increasing the number of agents, the convergence to a locally optimal policy takes longer (using one core of Intel Xeon Gold 6130), making it not too feasible for the larger setups we consider in this work. See Figure 8(a) for time taken to learn a SAC policy on a $1$-D cyclic topology with $\Delta t=3$, $N=M$, and same computational resources for all training setups of $N$. Our implementation of SAC was adapted from (Li, 2021). Furthermore, Figure 8(b) shows the performance of the learned SAC policy as compared to other algorithms. Although performance did improve as the number of agents rises, indicating the scalability of SAC to many agents, the performance of SAC remained suboptimal. Due to these limitations, we did not investigate the SAC algorithm further in our experiments.
@@
Figure 9. (a) Performance comparison for increased buffer capacity of each queue to $20$. The system utilization was increased to ensure occurrence of packet drops in a limited amount of time. (b) Performance comparison of MF-R when the servers are considered to be heterogeneous. Additional comparison was done with the Shortest-expected-delay algorithm (Selen et al., 2016), which is the state-of-the-art when the servers are heterogeneous.
D.4. Increased buffer size
@@ D.5. Heterogenous servers
We also conducted experiments where we considered the servers to be heterogeneous with randomly assigned speed of fast(rate $2$) or slow(rate $1$). The workload was the same; modulating between $[0.9,0.6]$. Fig. 9(b) shows that our algorithm has comparable performance to RND and OWN while outperforming JSQ. We also compared with Shortest-expected-delay (SED), the state-of-the-art load balancing algorithm for heterogeneous servers (Selen et al., 2016), which performs better since it makes decision based on both the neighbors’ queue state information and the server speeds. We believe our model is not an ideal fit for this setting and further work needs to be done in this direction.
Figure 10. (a): The training for TORUS with and without RNN policies converges almost to the same return. (b): Evaluation of the learned policy using different observations. MF-R uses only own queue state, MF-N additionally uses neighbour queue states, and MF-G uses state of all queues in the system.
D.6. RNNs and decentralized execution
We additionally trained our MF-R policy with recurrent neural networks in PPO, as is typical for partially-observed problems (Ni et al., 2022), MF-RNN. In Figure 10(a), we see that their effect is negligible and can even be negative. Finally, we used partially observed decentralized alternatives to the empirical distribution as input to the learned MF-R policy in the evaluation, see Figure 10(b). In MF-N we use the distribution of neighbours’ queue states, in MF-G we use the empirical distribution of all queues in the system, and in MF-R only the agent’s own queue state information is used as a one-hot vector. Similar performances indicate that the learned policy can be executed locally without using global information.